OpenStudy (anonymous):

Simplify the expression. 2/3x[15x5+(-24)+(-18)

OpenStudy (anonymous):

@Teddyiswatshecallsme @whpalmer4 @quickstudent @HeyWassup345 @Warga @JackofallTradez @automaticloveletter @iamabarbiegirl @Shadowless_Star @EclipsedStar

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

@sleepyjess @perl @khalilforthewin @jim_thompson5910

OpenStudy (anonymous):

@vera_ewing @freegirl11112 @geny03 @Hotchellerae21

OpenStudy (anonymous):

A 22 B48 C52 D16

OpenStudy (anonymous):

@confluxepic @TheSmartOne

OpenStudy (anonymous):

helppppppppppppppppp

OpenStudy (anonymous):



OpenStudy (anonymous):

@paigemc @tabdo1226 @Jessy02M @conad @Niguyver @beauty_queen67 @Preetha @HackberryAbby18 @TheOrion101

OpenStudy (texaschic101):

are the x's variables or are they just saying to multiply ??

OpenStudy (anonymous):

What he or she said

OpenStudy (texaschic101):

that would be " she "...lol

OpenStudy (anonymous):

ultiply

OpenStudy (anonymous):

What she said

OpenStudy (anonymous):

multiply

OpenStudy (texaschic101):

2/3 * [ 15 * 5 + (-24) + (-18) ] = 2/3 * [ 75 - 24 - 18 ] = 2/3 * 33 = 66/3 = 22

OpenStudy (anonymous):

What he or she said

OpenStudy (texaschic101):

questions ?

OpenStudy (anonymous):

nope thanks

OpenStudy (texaschic101):

anytime :)